# Bounds for the relative class number problem for function fields

### v1

@@ 1. Introduction
By contrast, the relative class number one problem for function fields was first considered by Leitzel–Madan [9] but is now fully resolved [5, 6, 7]. It is therefore natural to consider the problem of classifying finite separable extensions of function fields (of curves over finite fields) with relative class number equal to a fixed integer $m>1$. It is also natural to restrict to extensions which are either constant (i.e., induced by an extension of the constant field) or purely geometric (i.e., with no change in the constant field), as any extension can be interpreted as a constant extension followed by a purely geometric extension.
@@ 1.1. Bounds for the class number problem
Let $F^{\prime}/F$ be a finite separable extension of function fields of degree $d$ with relative class number $m$. Assume either that $F^{\prime}/F$ is constant and $d>1$, or that $F^{\prime}/F$ is purely geometric and $g^{\prime}>g$. Let $\mathbb{F}_{\_}q$ be the constant field of $F$. Let $g$ be the genus of $F$. Let $g^{\prime}$ be the genus of $F^{\prime}$.
@@ 1.2. The relative class number 2 problem
See Proposition 4.1 for an overview of the result and for further navigation into the proof (including tables of examples).
As in [5, 6, 7], the argument here mostly involves work over $\mathbb{F}_{\_}3$ and $\mathbb{F}_{\_}4$, including a number of exhaustive searches for cyclic covers of particular curves (using explicit class field theory as implemented in Magma) and a few arguments to rule out most noncyclic covers of degree 3 or 4. One advantage over the case $m=1$ is that we can easily rule out cyclic covers of prime degree greater than 2 (Lemma 2.9.) One disadvantage is that we do have to account for one noncyclic cubic cover; fortunately, we end up with enough circumstantial information about the cover that it is relatively straightforward to prove its existence and even to identify it explicitly (Lemma 7.3).
The case of the relative class number 2 problem in which the base field has constant field $\mathbb{F}_{\_}2$ is expected to be substantially harder due to the large genera of curves involved (although again there is no need to worry about cyclic covers of odd prime degree). Recall that in [5] one finds an example of a quadratic extension of relative class number 1 where the base field has genus 7, and the arguments in [7] to rule out further such extensions require a restricted census over the $\mathbb{F}_{\_}2$-points of the moduli space of curves of genus 7.
@@ 2. General facts
Let $A$ denote the Prym variety associated to the covering $C^{\prime}\to C$, i.e., the kernel of the map
@@
where $\operatorname{Res}$ denotes the Weil restriction. We then have $\dim(A)=g^{\prime}-g$ and #$\#A(\mathbb{F}_{\_}q)=h_{\_}{F^{\prime}/F}$.
@@ 2.1. Bounds on Prym varieties
The fundamental tool in studying the relative class number for function fields is the Weil bound for the Prym variety $A$, plus some refinements for small $q$.
@@ 2.1.1. $q>2$
Let $B$ be a simple abelian variety defined over $\mathbb{F}_{\_}q$ for $q>2$. Assume that $B$ is not in the isogeny classes 1.3.ad, 1.4.ae. Then
@@
When $q>4$, the result follows from the Weil bound. When $q=3,4$, the result follows from [4, Theorem 3.2] with the given exceptions. ∎
@@ 2.1.2. $q=2$
This approach cannot be directly adapted to $q=2$ because there are infinitely many simple abelian varieties of order 1 over $\mathbb{F}_{\_}2$ [10], so there is no hope of working with them concretely. Instead, we give a lower bound for a suitable linear combination of traces of an abelian variety depending on its order.
@@
By comparison, when #$\#A(\mathbb{F}_{\_}2)=1$, [5, Lemma 5.6] yields
@@ 2.2. Bounds on point counts
We will play bounds off against upper bounds on rational points derived from nonnegativity constraints.
@@
We adapt [5, Lemma 4.1] using the following values of $x_{\_}1$ and $x_{\_}2$.
@@
For explicit computations, we supplement with some upper bounds for the number of points #$\#C(\mathbb{F}_{\_}q)$ for a curve $C$ of genus $g$, taken from manYPoints.
@@
For $q=2$, in order to compare against Lemma 2.2, we use a bound including some point counts over extension fields.
@@
We use the parameters $x_{\_}2=0.8$ and $x_{\_}3=0.6$. See [5, Lemma 4.2]. The calculations can be found in the file linear-programming.ipybn. ∎
@@ 2.3. Constant case
Let $F^{\prime}/F^{\prime\prime}/F$ be a tower of proper constant extensions, with corresponding fields of constants $\mathbb{F}_{\_}{q^{\prime}}/\mathbb{F}_{\_}{q^{\prime\prime}}/\mathbb{F}_{\_}q$. Assume also that $g_{\_}F>0$ and that $h_{\_}{F^{\prime}/F}=2$. Then,
@@
[…]
@@
    If $h_{\_}{F^{\prime}/F^{\prime\prime}}=1$, then [5, Theorem 1.1] gives that $q^{\prime\prime}\leq 4$.
@@
    If $h_{\_}{F^{\prime}/F^{\prime\prime}}=2$, then $h_{\_}{F^{\prime\prime}/F}=1$ and [5, Theorem 1.1] gives that $q\leq 4$. ∎
@@ 2.4. Purely geometric case
When $F^{\prime}/F$ is purely geometric, we can rewrite Lemma 2.1 as an upper bound on the dimension in terms of $m=h_{\_}{F^{\prime}}/h_{\_}F$:
(2.7) $g^{\prime}-g=\dim(A)\leq\dfrac{\log(m)}{\log(b_{\_}q)}.$
A crucial observation that we will need is the following upper bound on the trace of the Prym variety [5, Equation (5.3)].
@@
$0\leq\#C^{\prime}(\mathbb{F}_{\_}{q^{r}})$ $=q^{r}+1-T_{\_}{C^{\prime},q^{r}},$
$=q^{r}+1-T_{\_}{C,q^{r}}-T_{\_}{A,q^{r}},$
=#$=\#C(\mathbb{F}_{\_}{q^{r}})-T_{\_}{A,q^{r}}.\qed$
Similarly, _$\sum_{\_}{n\mid r}\mu\mathopen{}\mathclose{{}\left(\frac{r}{n}}\right)\#C^{\prime}(\mathbb{F}_{\_}{q^{n}})$ equals $r$ times the number of degree-$r$ places of $C^{\prime}$, and hence is also nonnegative.
When $g\geq 2$ and $d=2$, from [5, (7.7)] and Equation 2.9 we have
@@
For cyclic extensions, we have an analogue of Lemma 2.5
@@
Let $\zeta$ be a primitive $\ell$-th root of unity. Form the étale cohomology groups $H^{1}(C),H^{1}(C^{\prime})$ with coefficients in $\mathbb{Q}_{\_}{\ell^{\prime}}$ for some prime $\ell^{\prime}$. The quotient $H^{1}(C^{\prime})/H^{1}(C)$, as a representation of the cyclic group $\mathrm{C}_{\_}\ell$, splits into eigenspaces for the $\ell-1$ nontrivial characters of $\mathrm{C}_{\_}\ell$. Since Frobenius commutes with the action of $\mathrm{C}_{\_}\ell$, it respects the eigenspace decomposition; choosing one eigenspace and taking the product of $1-\alpha$ as $\alpha$ runs over the Frobenius eigenvalues in that eigenspace yields an element of $\mathbb{Q}(\zeta)$ with norm $h_{\_}{F^{\prime}/F}$. This proves the claim. ∎
@@ 2.5. Noncyclic extensions of degree 3 and 4
For $d=3$, if $F^{\prime}/F$ is not cyclic, then $F^{\prime\prime}/F$ has Galois group $S_{\_}3$ and so contains a quadratic subextension $F^{\prime\prime\prime}/F$ (the quadratic resolvent of the original cubic extension). This extension may be either constant or purely geometric. However, if there exists a degree-1 place of $F$ which does not lift to a degree-1 place of $F^{\prime}$ (which is automatic if #$\#C^{\prime}(\mathbb{F}_{\_}q)<\#C(\mathbb{F}_{\_}q)$), then as in [6, Lemma 4.5], we see that the quadratic resolvent must be purely geometric. In this case, $F^{\prime\prime\prime}/F$ corresponds to a covering $C^{\prime\prime\prime}\to C$ whose Prym $A^{\prime}$ has the property that
@@
For $d=4$, one possibility is that $F^{\prime}/F$ admits an intermediate subfield (this includes the cyclic case). Excluding this option leaves three other possibilities. One is that $F^{\prime\prime}/F$ has Galois group $A_{\_}4$; by [6, Lemma 4.5], this case can be excluded by ensuring that some degree-1 place of $F$ lifts to a degree-4 place of $F^{\prime}$. In the other two cases, $F^{\prime\prime}/F$ has Galois group $S_{\_}4$ and again has a quadratic resolvent $F^{\prime\prime\prime}/F$ which can be either constant or purely geometric.
@@ 3.1.1. $q>4$
From the Weil lower bound (Equation 2.7), we deduce that
@@
When $g\leq 1$, we obtain the upper bound in (1). When $g>1$, Riemann–Hurwitz gives $g^{\prime}-g\geq(d-1)(g-1)$, from which we obtain (2). ∎
@@
From Equation 2.7 we obtain that
@@ 3.1.2. $q\in\mathopen{}\mathclose{{}\left\{3,4}\right\}$
From Equations 2.3, 2.3 and 2.2 we have that:
@@
If $g\leq 1$, then $g^{\prime}-1\leq g^{\prime}-g$ and we obtain (1). If $g>1$, then Riemann–Roch gives $(d-1)(g-1)\leq(g^{\prime}-g)$ and we obtain (2); to minimize $g^{\prime}$ we first take $d=2$ and then $g^{\prime}=2g-1$. ∎
@@ 3.1.3. $q=2$
If $g\leq 1$, then $g^{\prime}-1\leq g^{\prime}-g$ and we obtain (1). If $g>1$, then Riemann–Roch gives $(d-1)(g-1)\leq(g^{\prime}-g)$ and we obtain (2); to minimize $g^{\prime}$ we first take $d=2$ and then $g^{\prime}=2g-1$. ∎
Again, one can improve this bound by distinguishing the cases $d=2$ and $d>2$, using Equation 2.10 in the former case; for $m=1$ this is the content of [5, Lemma 10.1].
@@ 3.3. $q>4$
Let $B$ be an abelian variety over $\mathbb{F}_{\_}q$, with $q>4$, and let $m$ be a positive integer. Suppose that #$\#B(\mathbb{F}_{\_}{q^{d}})=m\cdot\#B(\mathbb{F}_{\_}q)$ for some $d\geq 2$. Then
$\mathopen{}\mathclose{{}\left(\frac{q^{d/2}-1}{q^{1/2}+1}}\right)^{2\dim B}\leq m.$
@@
$(q^{d/2}-1)^{2\dim(B)}$$\leq\#B(\mathbb{F}_{\_}{q^{d}}),$
$=m\cdot\#B(\mathbb{F}_{\_}q)\leq m(q^{1/2}+1)^{2\dim B},$
@@
We now apply Lemma 3.5 in the case when $B=J$ is the Jacobian of a curve $C/\mathbb{F}_{\_}q$, corresponding to a constant extension of function fields $F^{\prime}/F$ with relative class number $m$, so that

### v2

@@ 1. Introduction
By contrast, the relative class number one problem for function fields was first considered by Leitzel–Madan [11] but is now fully resolved [6, 7, 8]. It is therefore natural to consider the problem of classifying finite separable extensions of function fields (of curves over finite fields) with relative class number equal to a fixed integer $m>1$. (We exclude inseparable extensions because every purely inseparable extension of function fields has relative class number 1.) It is also natural to restrict to extensions which are either constant (i.e., induced by an extension of the constant field) or purely geometric (i.e., with no change in the constant field), as any extension can be interpreted as a constant extension followed by a purely geometric extension.
@@ 1.1. Bounds for the class number problem
Let $F^{\prime}/F$ be a finite separable extension of function fields of degree $d$ with relative class number $m$. Let $g$ be the genus of $F$ and let $g^{\prime}$ be the genus of $F^{\prime}$. Assume either that $F^{\prime}/F$ is constant and $d>1$, or that $F^{\prime}/F$ is purely geometric and $g^{\prime}>g$. Let $\mathbb{F}_{\_}q$ be the constant field of $F$. Let $g$ be the genus of $F$. Let $g^{\prime}$ be the genus of $F^{\prime}$.
@@ 1.2. The relative class number 2 problem
A more precise statement can be found in Proposition 4.1, which also includes a description of the extensions that occur in cases (b), (c), and (d); the proof of this statement occupies §5 (for $q>4$), §6 (for $q=4$), and §7 (for $q=3$). The extensions that occur in case (e) are listed in §8.
As in [6, 7, 8], the argument here mostly involves work over $\mathbb{F}_{\_}3$ and $\mathbb{F}_{\_}4$, including a number of exhaustive searches for cyclic covers of particular curves (using explicit class field theory as implemented in Magma) and a few arguments to rule out most noncyclic covers of degree 3 or 4. One advantage over the case $m=1$ is that we can easily rule out cyclic covers of prime degree greater than 2 (Lemma 2.9). One disadvantage is that we do have to account for one noncyclic cubic cover; fortunately, we end up with enough circumstantial information about the cover that it is relatively straightforward to prove its existence and even to identify it explicitly (Lemma 7.3).
The case of the relative class number 2 problem in which the base field has constant field $\mathbb{F}_{\_}2$ is expected to be substantially harder, due to the large genera of curves involved (although again there is no need to worry about cyclic covers of odd prime degree). To put this in context, recall that in [6] one finds an example of a quadratic extension of relative class number 1 where the base field has genus 7, and the arguments in [8] to rule out further such extensions require a restricted census over the $\mathbb{F}_{\_}2$-points of the moduli space of curves of genus 7. It is likely that a full resolution of the relative class number 2 problem in full would require at least a restricted census of curves of even larger genus over $\mathbb{F}_{\_}2$, and it is unclear whether such a census would be feasible.
@@ 2. General facts
Let $A$ denote the Prym variety associated to the covering $C^{\prime}\to C$, i.e., the reduced connected component of the identity of the kernel of the map
@@
where $\operatorname{Res}$ denotes the Weil restriction. We then have $\dim(A)=g^{\prime}-g$ and #$\#A(\mathbb{F}_{\_}q)=h_{\_}{F^{\prime}/F}$. Moreover,
(2.1) $h_{\_}{F^{\prime}/F}=P_{\_}A(1)=\prod_{\_}{\alpha}(1-\alpha),$
where the product ranges over the $q$-Frobenius eigenvalues of $A$. See [1, Section 2.6] for generalities on Prym schemes.
@@ 2.1. Bounds on Prym varieties
The fundamental tool in studying the relative class number for function fields is the Weil bound for the Prym variety $A$, plus some refinements for small $q$. See [5], and the references therein, for a general reference on this topic.
@@ 2.1.1. $q>2$
Let $B$ be an abelian variety over $\mathbb{F}_{\_}q$ for $q>2$ with no isogeny factor in the isogeny classes 1.3.ad, 1.4.ae. Then
@@
We immediately reduce to the case where $B$ is simple and not in one of the indicated isogeny classes. When $q>4$, the result follows from the Weil bound. When $q=3,4$, the result follows from [5, Theorem 3.2] with the given exceptions. ∎
@@ 2.1.2. $q=2$
This approach cannot be directly adapted to $q=2$ because there are infinitely many simple abelian varieties of order 1 over $\mathbb{F}_{\_}2$ [12], so there is no hope of working with them concretely. Instead, we give a lower bound for a suitable linear combination of traces of an abelian variety depending on its order. See [9] for a more detailed discussion.
@@
Figure 1. Positivity of the function defined by inequality (2.6).
By comparison, when #$\#A(\mathbb{F}_{\_}2)=1$, [6, Lemma 5.6] yields
@@ 2.2. Bounds on point counts
We will play bounds off against upper bounds on rational points derived from nonnegativity constraints. See [13, Chapter 6] for a general reference on this topic.
@@
We adapt [6, Lemma 4.1] using the following values of $x_{\_}1$ and $x_{\_}2$.
@@
For explicit computations, we supplement with some upper bounds for the number of points #$\#C(\mathbb{F}_{\_}q)$ for a curve $C$ of genus $g$, taken from manYPoints [15].
@@
The following lemma, when compared with Lemma 2.2, will give rise to bounds on the invariants of an extension $F^{\prime}/F$ in the case that $q=2$. See Lemma 3.4.
@@
We use the parameters $x_{\_}2=0.8$ and $x_{\_}3=0.6$. See [6, Lemma 4.2]. The calculations can be found in the file linear-programming.ipybn. ∎
@@ 2.3. Constant case
Let $F^{\prime}/F^{\prime\prime}/F$ be a tower of proper constant extensions, with corresponding fields of constants $\mathbb{F}_{\_}{q^{\prime}}/\mathbb{F}_{\_}{q^{\prime\prime}}/\mathbb{F}_{\_}q$. Assume also that $g_{\_}F>0$ and that $h_{\_}{F^{\prime}/F}=2$. Then:
@@
Recall that $h_{\_}{F^{\prime}/F}=h_{\_}{F^{\prime}/F^{\prime\prime}}\cdot h_{\_}{F^{\prime\prime}/F}$.
@@
    If $h_{\_}{F^{\prime}/F^{\prime\prime}}=1$, then [6, Theorem 1.1] gives that $q^{\prime\prime}\leq 4$.
@@
    If $h_{\_}{F^{\prime}/F^{\prime\prime}}=2$, then $h_{\_}{F^{\prime\prime}/F}=1$ and [6, Theorem 1.1] gives that $q\leq 4$. ∎
@@ 2.4. Purely geometric case
A crucial observation that we will need is the following upper bound on the trace of the Prym variety [6, Equation (5.3)].
@@
$0\leq\#C^{\prime}(\mathbb{F}_{\_}{q^{r}})$ $=q^{r}+1-T_{\_}{C^{\prime},q^{r}}$
$=q^{r}+1-T_{\_}{C,q^{r}}-T_{\_}{A,q^{r}}$
=#$=\#C(\mathbb{F}_{\_}{q^{r}})-T_{\_}{A,q^{r}}.$
Similarly, _$\sum_{\_}{n\mid r}\mu\mathopen{}\mathclose{{}\left(\frac{r}{n}}\right)\#C^{\prime}(\mathbb{F}_{\_}{q^{n}})$ equals $r$ times the number of degree-$r$ places of $C^{\prime}$, and hence is also nonnegative. That means that
00 0 _$\leq\sum_{\_}{n\mid r}\mu\mathopen{}\mathclose{{}\left(\frac{r}{n}}\right)(\#C(\mathbb{F}_{\_}{q^{n}})-T_{\_}{A,q^{n}})$
=_$=\sum_{\_}{n\mid r}\mu\mathopen{}\mathclose{{}\left(\frac{r}{n}}\right)\#C(\mathbb{F}_{\_}{q^{n}})-\sum_{\_}{n\mid r}\mu\mathopen{}\mathclose{{}\left(\frac{r}{n}}\right)T_{\_}{A,q^{n}}$
$=ra_{\_}r-\sum_{\_}{n\mid r}\mu\mathopen{}\mathclose{{}\left(\frac{r}{n}}\right)T_{\_}{A,q^{n}}.\qed$
When $g\geq 2$ and $d=2$, from [6, (7.7)] and Equation 2.10 we have
@@
For cyclic extensions, we have an analogue of Lemma 2.5.
@@
Let $\zeta$ be a primitive $\ell$-th root of unity. Form the étale cohomology groups $H^{1}(C)$ and $H^{1}(C^{\prime})$ with coefficients in $\overline{\mathbb{Q}_{\_}{\ell^{\prime}}}$ for some prime $\ell^{\prime}$ not equal to the characteristic of $F$. The quotient $H^{1}(C^{\prime})/H^{1}(C)$, as a representation of the cyclic group $\mathrm{C}_{\_}\ell$, splits into eigenspaces for the $\ell-1$ nontrivial characters of $\mathrm{C}_{\_}\ell$. Since Frobenius commutes with the action of $\mathrm{C}_{\_}\ell$, it respects the eigenspace decomposition; moreover, we may obtain all of the Frobenius eigenvalues appearing in $H^{1}(C^{\prime})/H^{1}(C)$ by taking the eigenvalues in one eigenspace together with their Galois conjugates over $\mathbb{Q}(\zeta)$. Consequently, the product of $1-\alpha$ as $\alpha$ runs over the Frobenius eigenvalues in one eigenspace is an element of $\mathbb{Q}(\zeta)$ with norm $h_{\_}{F^{\prime}/F}=P_{\_}A(1)$. This proves the claim. ∎
@@ 2.5. Noncyclic extensions of degree 3 and 4
For $d=3$, if $F^{\prime}/F$ is not cyclic, then $F^{\prime\prime}/F$ has Galois group $S_{\_}3$ and so contains a quadratic subextension $F^{\prime\prime\prime}/F$ (the quadratic resolvent of the original cubic extension). This extension may be either constant or purely geometric. However, if there exists a degree-1 place of $F$ which does not lift to a degree-1 place of $F^{\prime}$ (which is automatic if #$\#C^{\prime}(\mathbb{F}_{\_}q)<\#C(\mathbb{F}_{\_}q)$), then as in [7, Lemma 4.5], we see that the quadratic resolvent must be purely geometric. In this case, $F^{\prime\prime\prime}/F$ corresponds to a covering $C^{\prime\prime\prime}\to C$ whose Prym $A^{\prime}$ has the property that
@@
For $d=4$, one possibility is that $F^{\prime}/F$ admits an intermediate subfield (this includes the cyclic case). Excluding this option leaves three other possibilities. One is that $F^{\prime\prime}/F$ has Galois group $A_{\_}4$; by [7, Lemma 4.5], this case can be excluded by ensuring that some degree-1 place of $F$ lifts to a degree-4 place of $F^{\prime}$. In the other two cases, $F^{\prime\prime}/F$ has Galois group $S_{\_}4$ and again has a quadratic resolvent $F^{\prime\prime\prime}/F$ which can be either constant or purely geometric.
@@ 3.1.1. $q>4$
Since $F^{\prime}/F$ is purely geometric, we employ the exponential lower bounds on the number of points of Lemma 2.1, to get an upper bound on the difference $g^{\prime}-g$:
(3.1) $g^{\prime}-g=\dim(A)\leq\dfrac{\log(m)}{\log(b_{\_}q)}.$
From the Weil lower bound we take $b_{\_}q=(\sqrt{q}-1)^{2}$, and then we deduce that
@@
When $g\leq 1$, we obtain the upper bound in (1). When $g>1$, Riemann–Hurwitz gives $g^{\prime}-g\geq(d-1)(g-1)\geq(g-1)$, and therefore $g^{\prime}\geq 2g-1$ or equivalently $g\leq\frac{1}{2}(g^{\prime}+1)$. Using this bound on $g$ on
$g^{\prime}\leq\dfrac{\log(m)}{2\log(\sqrt{q}-1)}+g$
we obtain (2). Explicitly, we have
$g^{\prime}$ $\leq\dfrac{\log(m)}{2\log(\sqrt{q}-1)}+\frac{1}{2}(g^{\prime}+1),$
$2g^{\prime}$ $\leq\dfrac{\log(m)}{\log(\sqrt{q}-1)}+g^{\prime}+1,$
$g^{\prime}$ $\leq\dfrac{\log(m)}{\log(\sqrt{q}-1)}+1.$
@@
As in the proof of Lemma 3.1 we have
@@ 3.1.2. $q\in\mathopen{}\mathclose{{}\left\{3,4}\right\}$
From Equations 2.4, 2.3 and 2.3 we have that:
@@
If $g\leq 1$, then $g^{\prime}-1\leq g^{\prime}-g$ and we obtain (1). If $g>1$, then Riemann–Hurwitz gives $(d-1)(g-1)\leq(g^{\prime}-g)$ and we obtain (2); to minimize $g^{\prime}$ we first take $d=2$ and then $g^{\prime}=2g-1$. ∎
@@ 3.1.3. $q=2$
If $g\leq 1$, then $g^{\prime}-1\leq g^{\prime}-g$ and we obtain (1). If $g>1$, then Riemann–Hurwitz gives $(d-1)(g-1)\leq(g^{\prime}-g)$ and we obtain (2); to minimize $g^{\prime}$ we first take $d=2$ and then $g^{\prime}=2g-1$. ∎
Again, one can improve this bound by distinguishing the cases $d=2$ and $d>2$, using Equation 2.11 in the former case; for $m=1$ this is the content of [6, Lemma 10.1].
@@ 3.3. $q>4$
Let $J$ be an abelian variety over $\mathbb{F}_{\_}q$, with $q>4$, and let $m$ be a positive integer. Suppose that #$\#J(\mathbb{F}_{\_}{q^{d}})=m\cdot\#J(\mathbb{F}_{\_}q)$ for some $d\geq 2$. Then
$\mathopen{}\mathclose{{}\left(\frac{q^{d/2}-1}{q^{1/2}+1}}\right)^{2\dim(J)}\leq m.$
@@
$(q^{d/2}-1)^{2\dim(J)}$$\leq\#J(\mathbb{F}_{\_}{q^{d}})$
$=m\cdot\#J(\mathbb{F}_{\_}q)\leq m(q^{1/2}+1)^{2\dim(J)},$
@@
We now apply Lemma 3.5 in the case when $J$ is the Jacobian of a curve $C/\mathbb{F}_{\_}q$, corresponding to a constant extension of function fields $F^{\prime}/F$ with relative class number $m$, so that